# A Survey of Graph Unlearning

### v3

@@ 2 Preliminaries
Graph unlearning is an essential extension of the unlearning process to the domain of GML, where the input is an attributed graph $G$. Formally, let $\mathcal{M}$ be the original graph machine learning model, trained on the graph $G$ or a set of graphs $\mathcal{G}$ to perform a specific downstream task. The graph unlearning process aims to update $\mathcal{M}$ in the presence of data removal in the form of node or node feature vector, or edge or edge feature vector, etc. An illustration of graph unlearning is shown in Figure 2.
@@
Figure 2: The graph unlearning framework is illustrated. It depicts three types of removal requests, namely node removal, edge removal, and node feature removal. Each of these requests represents different scenarios where specific elements are to be removed from the trained model.
Definition 2.8 (Node Unlearning)
@@
Node unlearning aims to effectively remove the data associated with an individual node, denoted as $v$, from the trained model $\mathcal{M}$ and its corresponding training graph $G$. When a data subject $v$ exercises the right to revoke their data, node unlearning ensures the erasure of $v$’s node features, represented by $X_{v^{\prime}}$, as well as the links connecting $v$ with other nodes, denoted by $e_{u,v}$ for all $u$ in the neighborhood $\mathcal{N}_{v}$. Formally, node unlearning involves deriving an unlearned model $\mathcal{M}^{\prime}$ that is trained on the updated graph $G_{v}=G\setminus\{v,X_{v^{\prime}},e_{u,v}|\forall u\in\mathcal{N}_{v}\}$.
Definition 2.9 (Node Feature Unlearning)
@@
Node feature unlearning aims to obtain an unlearned model $\mathcal{M}^{\prime}$ that is trained on the modified graph $G_{v}=(V,E,X_{v^{\prime}},W)$, where $X_{v^{\prime}}$ represents the updated node feature matrix after removing the feature vector of node $v$. The removal from $X$ can be done by replacing the feature vector with zeros or ones. The model $\mathcal{M}^{\prime}$ is optimized to adapt to this updated graph $G_{v}$ while ensuring that the influence of $X_{v}$ on the model predictions is significantly reduced.
Definition 2.10 (Edge Unlearning)
@@
Edge unlearning refers to the removal of a specific edge from the trained model $\mathcal{M}$ and its corresponding training graph $G$, while retaining all nodes in the graph. In edge unlearning, a particular edge, denoted as $e_{u,v}$ connecting nodes $u$ and $v$, is requested to be revoked by the data subject or individual to whom it pertains. This process involves removing the edge along with its corresponding attributes, represented by $X_{e({u,v})}$, from the model’s training graph $G$, without affecting other connections and node features. Formally, edge unlearning seeks to obtain an unlearned model $\mathcal{M}^{\prime}$ that is trained on the updated graph $G^{\prime}=(V,E\setminus{(u,v)},X,X_{-e({u,v})})$, where $E_{-e_{(u,v)}}$ represents the graph with the edge removed $e_{(u,v)}$ and $X_{-e_{(u,v)}}$ corresponds to the updated edge feature matrix.
Unlearning in Graph-level Tasks
Unlearning in graph-level tasks where a set of graphs $\mathcal{G}=\{G_{1},G_{2},...,G_{N}\}$ is available, unlearning involves the removal of specific data points at different granularities. The graph-level tasks could be considered in a setting such as graph classification and graph regression. Mainly, the following types of unlearning requests can be received in this scenario:
  1. 1.
    Node unlearning request: A removal request for a specific node $v$ from a particular graph $G_{i}$ in the set can be received. This request requires updating the model parameters to accommodate the removal of node $v$ and its associated information from the graph $G_{i}$.
  2. 2.
    Set of nodes unlearning request: A removal request for a set of nodes $V^{\prime}=\{v_{i},v_{j},v_{k}\}$ from a subset of graphs $\{G_{i},G_{j},G_{k}\}$ can be received. In this scenario, the model parameters need to be updated to handle the removal of the specified nodes from the respective graphs without affecting the remaining graph data.
  3. 3.
    Edge or a set of edge unlearning request: A removal request for a specific edge $e_{(u,v)}$ from one graph or a set of edges that corresponds to multiple graphs from the set can be received.
  4. 4.
    Graph-level unlearning request: A request for the entire graph $G$ to be removed from $\mathcal{G}$ can also be received. This requires updating the model parameters to account for the removal of the entire graph $G$ and its associated information from the set of graphs $\mathcal{G}$.
@@ 3 Categorization and Frameworks
Drawing from the existing literature, we classify graph unlearning into two distinct categories: exact unlearning, and approximate unlearning. In Table 2, we offer a concise overview of these methods, providing a glimpse into their essence. Moving forward, in the forthcoming paragraphs, we will delve into each of these categories, providing explanations of the methodologies employed in graph unlearning. Our aim is to shed light on the intricacies of these methods, uncovering their unique contributions to the fascinating domain of graph unlearning.
Table 2: Categorization of graph unlearning approaches. $V\mathcal{M}^{\prime}$ denotes node unlearning, $E\mathcal{M}^{\prime}$ represents edge unlearning, $X\mathcal{M}^{\prime}$ refers to node feature unlearning, $G\mathcal{M}^{\prime}$ represents the graph unlearning approach, and graph uni. indicates graph universal methods. Approx. unl. stands for approximate unlearning, FL for federated learning, and ACM Tran. IST refers to ACM Transaction on Intelligent Systems and Technology. The Code column indicates the availability of the implementation for the proposed method.
Method Year&Venue Code $\boldsymbol{V\mathcal{M}^{\prime}}$ $\boldsymbol{E\mathcal{M}^{\prime}}$ $\boldsymbol{X\mathcal{M}^{\prime}}$ $\boldsymbol{G\mathcal{M}^{\prime}}$ $G$ Type $\mathcal{M}^{\prime}$ Type Method Summary
GraphEraser[17] SIGSAC2022 undirected exact unl. Cluster the graph, train a separate model on each cluster, and then aggregates the results
GUIDE [26] USENIX2023 undirected exact unl. Compute $k$ shards, apply repair function to recover edges, train $k$ models and aggregate
RecEraser[27] WWW2022 bipartite exact unl. SISA based framework to preserve the collaborative information in data partition
UltraRE[28] NeurIPS2024 bipartite exact unl. SISA based framework with balanced data partition and simplified combination process
GraphEditor[29] arXiv.2022 undirected exact unl. Consider non-convex setting, compute closed-from solution of GNN output and $\mathcal{Y}$ and update the model
@@
EraEdge[31] arXiv.2022 undirected approx. unl. Estimate the edge influence and propose EraEdge method to update the model
Certified Unl.[32] NeurIPS2022 undirected approx. unl. Use Hessian for updating the model and provides theoretical foundation
GST Unlearn[20] WWW2023 undirected approx. unl. Use the traditional influence functions to compute the change and update the model
@@
GNNDelete[21] ICLR2023 undirected approx. unl. Introduce shared weight matrices across the nodes and use layer-wise deletion operator to udpate the model
SGC[18] ICLR2023 undirected approx. unl. Use influence function for model updates and derive robust theoretical guarantees in convex setting
SAFE[19] arXiv2023 undirected approx. unl. Using graph sharding mechansim to train secure GNN models
FedLU[35] WWW2023 KG approx. unl. Federated Learning based learning and unlearning approach
SGNN[36] BigData2019 undirected FL Consider node feature similarity matrix and use FL to train a GNN
Feddy[37] ACM TIST2022 dynamic FL Perform secure aggregation and train GNN in FL setting
SCIF[39] Expert Syst. Appl. bipartite approx. unl. Selective and collaborative influence function in recommendation
IFRU[39] arxiv bipartite approx. unl. Influence function-based recommendation unlearing
@@ 3.1.1 SISA-based Retraining
A representative exact unlearning algorithm is an ensemble framework called SISA (Sharded, Isolated, Sliced, and Aggregated) [2]. Unlike traditional retraining which relies on the entire training dataset, SISA randomly partitions the training dataset into distinct and disjoint shards, each independently trained to form a set of shard models. Once a removal operation happens, only the relevant shard models need to be retrained. While SISA being the state-of-the-art solution for exact unlearning, directly applying SISA to graph domain might incure a destory of the graph structure, resulting in a significant drop in utility performance. To overcome this challenge, GraphEraser [17] is proposed as one of the pioneering works in graph unlearning. Similar to SISA, GraphEraser begins by partitioning the original training graph into separate and disjoint shards. Each shard is then used to train an individual model. When a node requires a prediction, GraphEraser forwards to all shard models and obtains their corresponding posteriors. These posteriors are then aggregated each shard model, enabling GraphEraser to make a robust prediction. Furthermore, when a node makes an unlearning request, GraphEraser promptly removes the node from the corresponding shard and subsequently retrains the shard model. This process ensures that the influence of this node is mitigated, and the model remains updated and accurate without requiring a complete retraining of the entire graph machine learning model. The difference to SISA mainly lies in the data partition and aggregation components. GraphEraser proposes two strategies to achieve balanced graph partition. One strategy is based on community detection, which aims to divide the graph into groups of nodes with dense connections internally and sparse connections between groups. Another strategy relies on embedding clustering, which considers both the node feature and graph structural information for the partitioning. Nodes with similar embeddings are assigned to the same shard. To further ensure a balanced partition, a Balanced Embedding k-means algorithm is devised. Additionally, considering that the shard models contribute differently to the final prediction, GraphEraser propose a learning-based aggregation method that optimizes the importance score of each shard model, which reflects its significance in the overall prediction process.
Following GraphEraser, which is a transductive unlearning approach, the authors in [26] present a method called GUIDE (GUided InDuctivE Graph Unlearning). GUIDE consists of three essential components: guided graph partitioning, graph repairing, and aggregation. GUIDE begins by partitioning the graph into shards using newly proposed clustering techniques. Subsequently, GUIDE employs subgraph repair methods to restore lost connections and integrates them into their respective shards. This step is crucial to reconstruct the graph after partitioning where some of the connections are lost. Finally, GUIDE applies an aggregation method to learn a similarity score for each shard. This similarity score is then utilized for making inferences during the unlearning process.
@@
In recommendation scenarios, the SISA framework has been tailored to handle the user-item bipartite graph [27]. To address challenges in data partitioning, RecEraser introduces three novel partitioning algorithms that preserve collaborative information, which is crucial for recommendation. Specifically, these three partitions are based on the similarities of users, items, and interactions. Additionally, the partition strategies are designed to achieve a balanced partition for the unlearning efficiency. Following RecEraser, [28] rethinks from an ensemble-based perspective and further refines RecEraser from the three potential losses respectively (i.e., redundancy, relevance, and combination). In detail, UltraRE [28] integrates transport weights in the clustering algorithm to optimize the data partition process, and simplifies the combination estimator with simple architecture.
@@ 3.1.2 Parameter-based Retraining
Another way to implement exact unlearning is to store the historical parameters of the model for further adjustment.
[40] propose an unlearning approach for training the model using mini-batch stochastic gradient descent (SGD). Throughout the training process, the model parameters are saved at each iteration. When deletion requests are received, retraining commences solely from the iteration when the deleted data first appeared. This strategy optimizes the unlearning process, allowing for targeted updates and minimizing redundant retraining efforts.
The authors in [29] introduce GraphEditor, a graph representation learning and unlearning approach that supports both node and edge deletion and addition. GraphEditor offers an exact unlearning solution without necessitating extensive retraining. The core principle behind GraphEditor lies in transforming the conventional GNN training problem into an alternative problem with a closed-form solution. Once the initial training is completed and the weight matrix $W$ is obtained, GraphEditor computes the closed-form solution, returning $W^{\prime}$, the updated weight matrix, and $S^{*}$, the inverse correlation matrix for unlearning. In the event of a deletion request, GraphEditor updates the model parameters $W^{\prime}$ by first removing the effect of the requested node/edge on the remaining nodes and then updating the model. This targeted approach ensures efficient and precise unlearning, making GraphEditor a useful tool for managing graph data and maintaining model integrity. However, a key limitation of the paper lies in its applicability, as the majority of GNNs employed are non-linear. Consequently, the current results may have limited relevance in such scenarios. Furthermore, envisioning the generalization of the current approach to more intricate network structures presents challenges, as it heavily relies on the linear characteristics of the problem.
In recommendation scenario, Untraining Alternating Least Squares (Untrain-ALS) [30] proposes to fine-tune based on the pretrained weights. They retrain the model with new data after the removal until convergence. They show that mathematically retraining with ALS equals minimizing the model loss on the remaining data after removal, making it an exact unlearning method.
3.2 Approximate Unlearning
Approximate unlearning aims to estimate the influence of unlearning data, and directly removes the influence through parameter manipulation. While facing the challenges related to unlearning completeness, these methods are generally more efficient. We discuss them in the convex and non-convex settings in the following.

### v4

@@ 2.3 Graph Unlearning
Graph unlearning is an essential extension of the unlearning process to the domain of GML, where the input is an attributed graph $G$. Formally, let $\theta^{0}$ be the weight of the original graph machine learning model, trained on the graph $G$ or a set of graphs $\mathcal{G}$ to perform a specific downstream task. The graph unlearning process aims to update $\theta^{0}$ in the presence of data removal in the form of node or node feature vector, or edge or edge feature vector, etc. An illustration of graph unlearning is shown in Figure 1.
@@
Figure 1: The graph unlearning framework is illustrated. Each removal request represents different scenarios where specific elements are to be removed from the trained model. Red color denotes the removal of component.
Transductive Learning. Graph machine learning when training on/for a single training graph $G$.
Definition 2.6 (Transductive Node Unlearning)
@@
Transductive node unlearning aims to effectively remove the data associated with an individual node, denoted as $v$, from the trained model $\theta^{0}$ and its corresponding training graph $G$. When a data subject $v$ exercises the right to revoke their data, node unlearning ensures the erasure of $v$’s node features, represented by $X_{\setminus v}$, as well as the links connecting $v$ with other nodes, denoted by $e_{u,v}$ for all $u$ in the neighborhood $\mathcal{N}_{v}$. Formally, node unlearning involves deriving an unlearned model $\mathcal{M}^{\prime}$ that is trained on the updated graph $G_{\setminus v}=G\setminus\{v,X_{\setminus v},e_{u,v}|\forall u\in\mathcal{N}_{v}\}$.
Definition 2.7 (Transductive Node Feature Unlearning)
@@
Transductive node feature unlearning aims to obtain an unlearned model $\theta^{r}$ that is trained on the modified graph $G_{\setminus v}=(V,E,X_{\setminus v},W)$, where $X_{\setminus v}$ represents the updated node feature matrix after removing the feature vector of node $v$. The removal from $X$ can be done by replacing the feature vector with zeros or ones. The model $\theta^{r}$ is optimized to adapt to this updated graph $G_{\setminus v}$ while ensuring that the influence of $X_{\setminus v}$ on the model predictions is significantly reduced.
Definition 2.8 (Transductive Edge Unlearning)
@@
Transductive edge unlearning refers to the removal of the influence of a specific edge from the trained model $\theta^{0}$ and its corresponding training graph $G$, while retaining all nodes in the graph. In edge unlearning, a particular edge, denoted as $e_{(u,v)}$ connecting nodes $u$ and $v$, is requested to be revoked by the data subject or individual to whom it pertains. This process involves removing the edge along with its corresponding attributes from the model’s training graph $G$, without affecting other connections and node features. Formally, edge unlearning seeks to obtain an unlearned model $\theta^{r}$ that is trained on the updated graph $G_{\setminus e_{(u,v)}}=(V,E\setminus\{e_{(u,v)}\},X,X_{\setminus e_{(u,v)}})$, where $X_{\setminus e_{(u,v)}}$ corresponds to the updated edge feature matrix.
Inductive Learning. Graph machine learning when seeking to generalize beyond a single graph to completely unseen nodes or even new graphs. The following types of unlearning requests can be received in this scenario:
Definition 2.9 (Inductive Node Unlearning)
Inductive node unlearning responds to the query to remove the data associated with a set of nodes $V^{-}=\{v_{i},\dots,v_{k}\}$ from their corresponding training graphs $\mathcal{G}^{\prime}=\{G_{i},\dots,G_{k}\}$ in the dataset. Similar to the transductive equivalent, inductive node unlearning ensures the erasure of $v_{i}$’s node features, and the links connecting $v_{i}$ with adjacent nodes within graph $G_{i}$. Denoting $G^{\prime}_{i}$ as the altered graph $G_{i}$ with node $v_{i}$ removed as described, we can restate the new version formally: inductive node unlearning involves deriving an unlearned model $\mathcal{M}^{\prime}$ that is trained on the updated dataset $\mathcal{G}_{\setminus V^{-}}=\mathcal{G}\setminus\mathcal{G}^{\prime}\;\cup\;\{G^{\prime}_{i},\dots G^{\prime}_{k}\}$.
Definition 2.10 (Inductive Edge Unlearning)
Straightforwardly similar to Definition 2.9, inductive edge unlearning responds to the query to remove the data associated with a set of edges $E^{-}=\{e_{(u_{i},v_{i})},\dots,e_{(u_{k},v_{k})}\}$ from their corresponding training graphs $\mathcal{G}^{\prime}=\{G_{i},\dots,G_{k}\}$ in the dataset. Let $G^{\prime}_{i}$ be the altered graph $G_{i}$ with edge $e_{(u_{i},v_{i})}$ removed as specified in Definition 2.8, inductive edge unlearning involves deriving an unlearned model $\mathcal{M}^{\prime}$ that is trained on the updated dataset $\mathcal{G}_{\setminus E^{-}}=\mathcal{G}\setminus\mathcal{G}^{\prime}\;\cup\;\{G^{\prime}_{i},\dots G^{\prime}_{k}\}$.
Definition 2.11 (Graph-Level Unlearning)
This is a special case to the inductive setting, where the training dataset has more than one graph. A request to unlearn a set of graphs $\mathcal{G}^{-}=\{G_{i}^{-},\dots,G_{k}^{-}\}$ requires updating the model parameters to account for the removal of every graph in $\mathcal{G}^{-}$ and its associated information from the set of graphs $\mathcal{G}$. Formally, graph-level unlearning involves deriving an unlearned model $\mathcal{M}^{\prime}$ that is trained on the updated dataset $\mathcal{G}_{\setminus G^{-}}=\mathcal{G}\setminus\mathcal{G}^{-}$.
@@ 3 Categorization and Frameworks
Drawing from the existing literature, we classify graph unlearning into two distinct categories: exact unlearning, and approximate unlearning. In Table 2, we offer a concise overview of these methods, providing a glimpse into their essence. In the forthcoming paragraphs, we will delve into each of these categories, providing explanations of the methodologies employed in graph unlearning. Our aim is to shed light on the intricacies of these methods, uncovering their unique contributions to the fascinating domain of graph unlearning.
Table 2: Categorization of graph unlearning approaches. $V$ denotes node unlearning, $E$ represents edge unlearning, $X$ refers to node feature unlearning, $G$ represents the graph unlearning approach. Approx. unl. stands for approximate unlearning, Code indicates the availability of the implementation for the proposed method.
Method Year & Venue Code $\boldsymbol{V}$ $\boldsymbol{E}$ $\boldsymbol{X}$ $\boldsymbol{G}$ $G$ Type $\mathcal{M}$ Type Method Summary
GraphEditor [26] arXiv 2022 undirected exact unl. Compute closed-from solution of GNN output and $\mathcal{Y}$ and update the model for non-convex setting
RecEraser [27] WWW 2022 bipartite exact unl. SISA-based framework to preserve the collaborative information in data partition
@@
UltraRE [29] NeurIPS 2024 bipartite exact unl. SISA-based framework with balanced data partition and simplified combination process
GraphRevoker [30] WWW 2024 undirected exact unl. SISA-based method with graph-aware clustering and contrastive submodel aggregation.
SCIF [31] ESWA 2023 bipartite approx. unl. Selective and collaborative influence function in recommendation
IFRU [33] ACM ToRS 2024 bipartite approx. unl. Influence function-based recommendation unlearing
GraphEraser [15] SIGSAC 2022 undirected approx. unl. Cluster the graph, train a separate model on each cluster, and then aggregates the results
CGU [34] NeurIPS 2022 undirected approx. unl. Use Hessian for updating the model and provides theoretical foundation
GUIDE [35] USENIX 2023 undirected approx. unl. Compute $k$ shards, apply repair function to recover edges, train $k$ models and aggregate
CEU [36] KDD 2023 undirected approx. unl. Propose a custom edge influence function and Newton-update accordingly
Unlearning GST [14] WWW 2023 undirected approx. unl. Use the traditional influence functions to compute the change and update the model
@@
GNNDelete [16] ICLR 2023 undirected approx. unl. Introduce shared weight matrices across nodes; apply layer-wise deletion operator to update the model
SGC [39] ICLR 2023 undirected approx. unl. Use influence function for model updates, derive robust theoretical guarantees in convex setting
SAFE [13] ICCV 2023 undirected approx. unl. Using graph sharding mechanism to train secure GNN models
GUKD [40] ICICS 2023 undirected approx. unl. Finetune the unlearning model with soft labels from a smaller independently retrained model.
MEGU [41] AAAI 2024 undirected approx. unl. Simultaneously unlearn and preserve original utility through a set of high influence nodes
D2DGN [42] arXiv 2024 undirected approx. unl. Distill model to destroy knowledge of the unlearned data and retain knowledge of the remaining data
IDEA [43] KDD 2024 undirected approx. unl. Provides a certification for edge, node, and feature unlearning using influence function
UtU [44] WWW 2024 undirected approx. unl. Prevent message passing by removing unlearned edges during inference
Cognac [45] ICML 2025 undirected approx. unl. Detect nodes affected by the unlearning requests and apply contrastive unlearning on them
ETR [46] AAAI 2025 undirected approx. unl. Edit model weights using Fisher information matrix, then one-step gradient update for remaining data
ScaleGUN [47] ICLR 2025 undirected approx. unl. Maintain certification for large graphs by lazy local propagation and only retrain when exceeding tolerance threshold
KGU Schema [48] COLING 2025 undirected approx. unl. Query subgraphs similar to the removal request and unlearn them
AGU [49] IJCAI 2025 undirected approx. unl. Address edge removal and feature removal separately by identifying and correcting affected nodes
SGU [50] arXiv 2025 undirected approx. unl. Scalable GU through a set of high influence nodes with Node Influence Maximization
@@ 3.1.1 SISA-based Retraining
A representative exact unlearning algorithm is an ensemble framework called SISA (Sharded, Isolated, Sliced, and Aggregated) [2]. Unlike traditional retraining which relies on the entire training dataset, SISA randomly partitions the training dataset into distinct and disjoint shards, each independently trained to form a set of shard models. Once a removal operation happens, only the relevant shard models need to be retrained. In recommendation scenarios, the SISA framework has been tailored to handle the user-item bipartite graph [27]. To address challenges in data partitioning, RecEraser introduces three novel partitioning algorithms that preserve collaborative information, which is crucial for recommendation. Specifically, these three partitions are based on the similarities of users, items, and interactions. Additionally, the partition strategies are designed to achieve a balanced partition for the unlearning efficiency. Following RecEraser, [29] rethinks from an ensemble-based perspective and further refines RecEraser from the three potential losses respectively (i.e., redundancy, relevance, and combination). In detail, UltraRE [29] integrates transport weights in the clustering algorithm to optimize the data partition process, and simplifies the combination estimator with simple architecture. GraphRevoker [30] took a different approach to prevent information loss by proposing a different clustering algorithm named graph-aware sharding, and improving upon submodel ensembling with contrastive submodel aggregation.
However, note that some SISA-style methods contain modifications that made them no longer exact. Specifically, if the graph partitioning process is not fully random and instead clustered conditionally on edge or feature information, it voids the exactness property of the overall method. One example is GraphEraser [15] which begins by partitioning the original training graph into separate and disjoint shards. Each shard is then used to train an individual model. When a node requires a prediction, GraphEraser forwards to all shard models and obtains their corresponding posteriors. These posteriors are then aggregated each shard model, enabling GraphEraser to make a robust prediction. Furthermore, when a node makes an unlearning request, GraphEraser promptly removes the node from the corresponding shard and subsequently retrains the shard model. This process ensures that the influence of this node is mitigated, and the model remains updated and accurate without requiring a complete retraining of the entire graph machine learning model. The difference to SISA mainly lies in the data partition and aggregation components. GraphEraser proposes two strategies to achieve balanced graph partition. One strategy is based on community detection, which aims to divide the graph into groups of nodes with dense connections internally and sparse connections between groups. Another strategy relies on embedding clustering, which considers both the node feature and graph structural information for the partitioning. Nodes with similar embeddings are assigned to the same shard. To further ensure a balanced partition, a Balanced Embedding k-means algorithm is devised. Additionally, considering that the shard models contribute differently to the final prediction, GraphEraser propose a learning-based aggregation method that optimizes the importance score of each shard model, which reflects its significance in the overall prediction process.
Inspired by GraphEraser, Wang et al. present a method named GUIDE (GUided InDuctivE Graph Unlearning) [35]. GUIDE consists of three essential components: guided graph partitioning, graph repairing, and aggregation. GUIDE begins by partitioning the graph into shards using newly proposed clustering techniques. Subsequently, GUIDE employs subgraph repair methods to restore lost connections and integrates them into their respective shards. This step is crucial to reconstruct the graph after partitioning where some of the connections are lost. Finally, GUIDE applies an aggregation method to learn a similarity score for each shard. This similarity score is then utilized for making inferences during the unlearning process.
@@ 3.1.2 Parameter-based Retraining
Another way to implement exact unlearning is to store the historical parameters of the model for further adjustment. [51] propose an unlearning approach for training the model using mini-batch stochastic gradient descent (SGD). Throughout the training process, the model parameters are saved at each iteration. When deletion requests are received, retraining commences solely from the iteration when the deleted data first appeared. This strategy optimizes the unlearning process, allowing for targeted updates and minimizing redundant retraining efforts. The authors in [26] introduce GraphEditor, a graph representation learning and unlearning approach that supports both node and edge deletion and addition. GraphEditor offers an exact unlearning solution without necessitating extensive retraining. The core principle behind GraphEditor lies in transforming the conventional GNN training problem into an alternative problem with a closed-form solution. Once the initial training is completed and the weight matrix $W$ is obtained, GraphEditor computes the closed-form solution, returning $W^{\prime}$, the updated weight matrix, and $S^{*}$, the inverse correlation matrix for unlearning. In the event of a deletion request, GraphEditor updates the model parameters $W^{\prime}$ by first removing the effect of the requested node/edge on the remaining nodes and then updating the model. This targeted approach ensures efficient and precise unlearning, making GraphEditor a useful tool for managing graph data and maintaining model integrity. However, a key limitation of the paper lies in its applicability, as the majority of GNNs employed are non-linear. Consequently, the current results may have limited relevance in such scenarios. Furthermore, envisioning the generalization of the current approach to more intricate network structures presents challenges, as it heavily relies on the linear characteristics of the problem. In recommendation scenario, Untraining Alternating Least Squares (Untrain-ALS) [28] proposes to fine-tune based on the pretrained weights. They retrain the model with new data after the removal until convergence. They show that mathematically retraining with ALS equals minimizing the model loss on the remaining data after removal, making it an exact unlearning method.
3.2 Approximate Graph Unlearning
Approximate unlearning usually aims to estimate the influence of unlearning data, and directly removes the influence through parameter manipulation. While facing the challenges related to unlearning completeness, these methods are generally more efficient. We discuss them in the convex and non-convex settings in the following.